OpenStudy (anonymous):

What is the slope of y=6

undeadknight26 (undeadknight26):

6? not sure...

OpenStudy (solomonzelman):

ZERO as any horizontal line.

undeadknight26 (undeadknight26):

i agree with any thing that guy says.

OpenStudy (charlotte123):

Zero Would Be Correct! http://www.wolframalpha.com/input/?i=y%3D6+slope

OpenStudy (charlotte123):

@undeadknight26 LOL!

OpenStudy (solomonzelman):

vertical line y=c lets choose any points on line c (c,a) and (c,b) and plug them into the slope formula. \[\huge\color{blue}{\frac{c-c}{a-b}=\frac{0}{a-b}}\]